# Inertial range of magnetorotational turbulence

### v1

@@
Yohei Kawazura1,2∗, Shigeo S. Kimura1,3
@@ 6-3 Aoba, Aramaki, Sendai 980-8578, Japan
3Astronomical Institute, Tohoku University
@@
To whom correspondence should be addressed; E-mail: kawazura@tohoku.ac.jp
@@ Introduction
Recently, these two mysteries in magnetorotational turbulence, namely, the energy spectra and the partition of MHD modes, were resolved using the reduced MHD model  (?), which assumes the presence of a near-azimuthal mean magnetic field. More specifically, both the kinetic and magnetic spectra approach $k^{-3/2}$ with the same amplitude as the cascade proceeds where $k$ is the wavenumber, and the energy flux of the slow magnetosonic fluctuations is almost double that of the Alfvénic fluctuations. This study aims to resolve the inertial range of magnetorotational turbulence by leveraging the power of the world’s fastest supercomputer, and to examine whether the predictions made by the reduced MHD are valid or not.
@@
Here, we present the direct numerical simulation of magnetorotational turbulence with the highest resolution in history. The simulation was performed on the Fugaku supercomputer (the world’s fastest machine until May 2022), using approximately 128 million central processing unit (CPU) hours. We solved the incompressible MHD equations using a pseudospectral method in a local shearing box with $N_{x}\times N_{y}\times N_{z}=8192\times 8192\times 4096$ grid points, where $x,y$, and $z$ denote radial, azimuthal, and vertical directions, respectively. The initial magnetic field was set to be vertically uniform with small-amplitude perturbations. The size of the box was set to $L_{x}\times L_{y}\times L_{z}=4\lambda\times 8\lambda\times 2\lambda$, where $\lambda=2\pi v_{\mathrm{A}}/\Omega$ approximately equals the wavelength of the fastest-growing modes of MRI, $v_{\mathrm{A}}$ is the Alfvén speed given by the initial magnetic field, and $\Omega$ is the angular velocity of the accretion disk. To narrow the dissipation range, the cascade was terminated by fourth-order hyperviscosity and hyperresistivity. Further details of the numerical setup can be found in the Materials and Methods.
@@
Figure 1, A and B show the snapshot of the norm of the flow field $\mathbf{u}$ and the magnetic field $\mathbf{B}$ together with typical magnetic field lines, in the planes $x=0$, $y=0$, and $z=0$. Each field is normalized by its root mean square value. The overall structures of both fields are azimuthally elongated due to the Keplerian shear flow. Upon closer inspection, the magnetic field has broad structures, while the flow field is concentrated in relatively smaller patches. The reason for this difference is that the large-scale azimuthal magnetic field $B_{y}$ is preferentially amplified by the $\Omega$ effect  (?), which is evident from the shape of magnetic field lines on the $x=0$ and $z=0$ planes (see also Fig. S1 for the snapshot of all components of $\mathbf{u}$ and $\mathbf{B}$ showing that only $B_{y}$ has intense large-scale structures). One also notices that the magnetic field lines are predominantly azimuthal but have a finite radial component with anticorrelation between $B_{x}$ and $B_{y}$ (consistent with the hybrid PIC simulation of the shearing box magnetorotational turbulence  (?); see also Fig. S1, A and B). Figure 1, C and D show the same snapshot of $\mathbf{u}$ and $\mathbf{B}$ but high-pass filtered by removing the fluctuation with a wavenumber smaller than 20. As we will show in Fig. 2C and 3 below, the fluctuations with the wavenumber greater than $10$ are arguably in the inertial range. The filtered $\mathbf{u}$ and $\mathbf{B}$ exhibit similar morphology, unlike their unfiltered counterparts, suggesting that the cascade tends to be Alfvénic as it moves towards smaller scales. It is also observed that the spatial structures of the filtered fields are significantly elongated along the magnetic field lines. The magnified regions in Fig 1, C and D clearly visualize this elongation. Thus, the large-scale magnetic field effectively acts like a mean field for the fluctuations in the inertial range. One also finds that the azimuthal polarity of the mean magnetic field is not uniform, which is evident from the typical magnetic field lines plotted in Fig. 1. This is because MRI creates both positive and negative $B_{x}$, which turn into positive and negative $B_{y}$ via shear flow. Therefore, the emergence of the current sheet and magnetic reconnection of a mean field are naturally expected to occur in the simulation, and indeed, we find the chain of multiple plasmoids in the magnified region on the $x=0$ plane in Fig. 1C. Nevertheless, these plasmoids do not fill a significant portion of the simulation domain but appear only in limited regions.
@@
To explore the properties of small-scale fluctuations, we analyzed the energy spectra of $\mathbf{u}$ and $\mathbf{B}$. Figure 2A shows the omnidirectional spectra of kinetic, magnetic, and total energy. The total energy spectrum is well fitted by a power law of $k^{-5/3}$ when $k\lesssim 100$, where $k$ is the wavenumber, and the spectral slope gets slightly shallower when $k\gtrsim 100$. While the $-5/3$ spectrum was captured in the previously highest-resolution simulation  (?), the shallowing of the slope is found in this study for the first time. We found that the shallowing occurs only in the magnetic field, while the kinetic spectrum is almost completely $k^{-3/2}$ throughout the wavenumber domain. In the previously highest-resolution simulation  (?), the kinetic spectrum was slightly shallower than $k^{-3/2}$. This is likely due to insufficient numerical resolution because, as we will show in Fig. 4A (inset), the kinetic spectrum becomes shallower than $k^{-3/2}$ in our low-resolution simulation.
@@
We further investigate the magnetic spectrum by plotting $x,y$, and $z$ components in Fig. 2B. As is consistent with the previous simulation  (?), $B_{y}$ is dominant on the large scale (due to amplification via the $\Omega$ effect) and diminishes quickly in the smaller scales. However, unlike the previous simulation, the $B_{y}$ spectrum becomes shallower at $k\gtrsim 100$, at which the total energy spectrum starts to become shallower than $-5/3$. The second dominant component on the large scale is $B_{x}$, which is presumably due to MRI. Although one might doubt that the shallowing of the spectrum is the numerical roll-up due to the hyper-dissipation, we do not think that this is the case because of the following two reasons. First, only the $B_{y}$ spectrum shows shallowing, while the other fields do not. Second, as we will show in Fig. 4A (inset), the shallowing disappears in the low-numerical-resolution simulation. We will further investigate this flattening of $B_{y}$ later in this paper.
@@
Evidence that our simulation resolved the inertial range is provided in Fig. 2C, which plots the cross-scale energy flux through the wavenumber shell $|\mathbf{k}|=k$. Here, the energy flux $\Pi^{f^{<}}_{g^{>}}(k)$ denotes the energy transfer from the field $f$, where $f=u$ (or $=B$) for the flow (or magnetic) field, with the wavenumber smaller than $k$ to the field $g$ with the wavenumber larger than $k$ (see Materials and Methods for the mathematical definition). We find that the total energy flux is fairly constant at $k\gtrsim 10$, which means that the cascade at $k\gtrsim 10$ is in the inertial range. We stress that, in the previous simulations of MRI turbulence, the cross-scale energy flux was not constant  (?), and our simulations is the first time that found the constant energy flux. Note, however, that not only the number of grid points, but the size of the simulation box and the aspect ratio of their simulation are different from ours. We also find that the amount of energy flux “from $u$ and $B$ to $B$” dominates that “from $u$ and $B$ to $u$”, which is consistent with the observation by Ref.  (?).
@@
The nonlinear energy transfer is further investigated in Fig. 3. We define the transfer function $\mathcal{T}_{fg}(Q,K)$ that denotes the energy transfer from the field $f$ in the wavenumber shell $Q\leq|\mathbf{k}|<Q+1$ to the field $g$ in the wavenumber shell $K\leq|\mathbf{k}|<K+1$  (?, ?, ?, ?) (see Materials and Methods for the mathematical definition). We find that the dominant energy transfer among all possible combinations of fields is $\mathcal{T}_{BB}$, which is consistent with Ref.  (?). Noticeably, there is a transition at $|\mathbf{k}|\approx 4$. In $|\mathbf{k}|>4$, both $\mathcal{T}_{uu}$ and $\mathcal{T}_{BB}$ are fairly local, and the direction of the cascade is forward. On the other hand, in $|\mathbf{k}|<4$, none of the energy transfers are local. We also find that there is an inverse energy transfer for $\mathcal{T}_{BB}$, suggesting that the large-scale structures of the magnetic field shown in Fig. 1A is formed not just by the $\Omega$ effect but also by the inverse cascade. Furthermore, the transition scale, $|\mathbf{k}|\approx 4$, coincides with the scale where the energy flux becomes constant (Fig. 2C), meaning that the energy cascade in the inertial range is local. Regarding the energy transfers between $u$ and $B$, we find that they are more non-local than the transfer within $u$ or $B$ as $\mathcal{T}_{uB}$ and $\mathcal{T}_{Bu}$ have broad off-diagonal tails. That being said, the non-local transfer from the injection range ($|\mathbf{k}|\leq 4$) peters out as the cascade proceeds. This is evident from Fig. 3, E-H, which shows the contribution of energy transfer from the injection range defined by $\sum_{Q\leq 4}\mathcal{T}_{fg}(K,Q)/\sum_{Q\leq K-1}\mathcal{T}_{fg}(K,Q)$. We find that the contribution of the injection range in $\mathcal{T}_{uu}$ and $\mathcal{T}_{BB}$ disappears immediately below the transition scale $K=4$, which is obvious since these transfers are fairly local, as mentioned above. For $\mathcal{T}_{uB}$ and $\mathcal{T}_{Bu}$, the contribution of the injection range survives down to the relatively smaller scale, but the contribution becomes less than 10% at $K=50$ for $\mathcal{T}_{uB}$ and at $K=40$ for $\mathcal{T}_{Bu}$. This is dramatically different from L11, which showed that the box-scale to grid-scale transfer of $\mathcal{T}_{uB}$ was significant. In short, the two facts that the cross-scale energy flux is constant and that the non-local energy transfer from the injection range peters out manifest that the cascade in our simulation is in the inertial range. This is one of the two main results of this paper.
@@
The other main result of this paper, namely, the partition between slow-magnetosonic and Alfvénic fluctuations in the magnetorotational turbulence, is shown in Fig. 4. In incompressible MHD, the eigenfunctions of the slow-magnetosonic waves (also known as pseudo-Alfvén waves in the incompressible limit) and Alfvén waves are simply given by the parallel and perpendicular components of $\mathbf{u}$ and $\mathbf{B}$ to the mean magnetic field, which are denoted by $(u_{\|},\delta B_{\|})$ and $(u_{\perp},\delta B_{\perp})$, respectively. However, the global mean magnetic field does not always serve as a mean field for small-scale fluctuations, and thus, we use the method developed by Cho and Lazarian  (?) to decompose the fluctuation and the local mean magnetic field $\mathbf{B}_{0}(\mathbf{r})$. For a given wavenumber $k$, the local mean magnetic field $\mathbf{B}_{0}(\mathbf{r})$ is obtained by filtering the Fourier modes of $\mathbf{B}$ with the wavenumber greater than $k/2$, and the fluctuations are obtained by filtering out the Fourier modes of $\mathbf{u}$ and $\mathbf{B}$ with the wavenumber smaller than $k/2$ or greater than $2k$. Then we decompose the fluctuations of $\mathbf{u}$ and $\mathbf{B}$ into parallel ($u_{\|}$ and $\delta B_{\|}$) and perpendicular ($u_{\perp}$ and $\delta B_{\perp}$) components to $\mathbf{B}_{0}$, which correspond to slow-magnetosonic and Alfvénic fluctuations, respectively. Figure 4A shows the spectra of the decomposed fields. First, we confirm the validity of our decomposition by comparing the sum of $u_{\|}$ and $u_{\perp}$ with the total $u$ in Fig. 1A and the sum of $\delta B_{\|}$ and $\delta B_{\perp}$ with the total $B$ in Fig. 1A; the spectra of the total fields are almost perfectly recovered from the sum of the decomposed fields. In terms of magnetic fluctuations, one finds that only the spectrum of $\delta B_{\|}$ exhibits flattening at $k\gtrsim 100$, while $\delta B_{\perp}$ has nearly the same spectral index throughout the inertial range. In contrast, the spectra of both $u_{\|}$ and $u_{\perp}$ are almost perfectly $k^{-3/2}$. Therefore, the flattening of the spectrum seen in Fig. 1A is solely due to the magnetic component of slow-magnetosonic fluctuations. The inset of Fig. 4A shows the same spectra obtained from the simulation with lower resolution $N_{x}\times N_{y}\times N_{z}=512\times 512\times 256$, which manifests that the flattening of $\delta B_{\|}$ is absent. Thus, the flattening can be seen only with ultra-high resolution. It can be seen that due to this flattening, the magnetic energy and kinetic energy of slow-magnetosonic fluctuations tend to be equipartitioned, and both spectra approach $k^{-3/2}$. On the other hand, the kinetic and magnetic spectra of Alfvénic fluctuations do not converge in our simulation, and a further higher resolution is required to determine the converged spectral slope. Figure 4B shows the ratio of the energy of slow-magnetosonic fluctuations to that of Alfvénic fluctuations, manifesting that slow-magnetosonic fluctuations have approximately twice stronger energy than the Alfvénic ones. One also finds that the ratio is almost constant throughout the wavenumber domain, suggesting that that the coupling between the Alfvénic and slow-magnetosonic fluctuations in MRI turbulence is weak. This is consistent with the recent report on the shearing box simulation which found that the ratio of energy injection between Alfvénic and slow-magnetosonic fluctuations almost equals to the ratio of dissipation which we computed using the redueced MHD  (?). All the features mentioned above are consistent with the numerical simulation of magnetorotational turbulence solved by the reduced MHD with near-azimuthal mean magnetic field  (?); Figure 4C is the spectra of the slow-magnetosonic and Alfvénic fluctuations, and they are remarkably similar with the spectra at $k\gtrsim 100$ in Fig. 4A; Figure 4D is the ratio between the two fluctuations, and the value is $\approx 2$, which is also exactly the same as Fig. 4B.

### v2

@@
Yohei Kawazura1,2,3∗, Shigeo S. Kimura1,4
@@ 6-3 Aoba, Aramaki, Sendai 980-8578, Japan
3School of Data Science and Management, Utsunomiya University
350 Minemachi, Utsunomiya, Tochigi 321-8505 Japan
4Astronomical Institute, Tohoku University
@@
To whom correspondence should be addressed; E-mail: kawazura@a.utsunomiya-u.ac.jp
@@ Introduction
Recently, these two mysteries in magnetorotational turbulence, namely, the energy spectra and the partition of MHD modes, were resolved using the reduced MHD model  (?), which assumes the presence of a near-azimuthal mean magnetic field. More specifically, both the kinetic- and magnetic-energy spectra approach $k^{-3/2}$ with the same amplitude as the cascade proceeds where $k$ is the wavenumber, and the energy flux of the slow magnetosonic fluctuations is almost double that of the Alfvénic fluctuations. This study aims to resolve the inertial range of magnetorotational turbulence by leveraging the power of the world’s fastest supercomputer, and to examine whether the predictions made by the reduced MHD are valid or not.
Results
Here, we present the direct numerical simulation of magnetorotational turbulence with the highest resolution in history. The simulation was performed on the Fugaku supercomputer (the world’s fastest machine until May 2022), using approximately 128 million central processing unit (CPU) hours. We solved the incompressible MHD equations using a pseudospectral method in a local shearing box with $N_{x}\times N_{y}\times N_{z}=8192\times 8192\times 4096$ grid points, where $x,y$, and $z$ denote radial, azimuthal, and vertical directions, respectively. The initial magnetic field was set to be vertically uniform with small-amplitude perturbations. The size of the box was set to $L_{x}\times L_{y}\times L_{z}=4\lambda\times 8\lambda\times 2\lambda$, where $\lambda=2\pi v_{\mathrm{A}}/\Omega$ approximately equals the wavelength of the fastest-growing modes of MRI, $v_{\mathrm{A}}$ is the Alfvén speed given by the initial magnetic field, and $\Omega$ is the angular velocity of the accretion disk. We note that the size of our simulation box is relatively smaller than those used in other simulations. This configuration was chosen because our objective is to investigate the inertial range, which is expected to appear at scales smaller than $\lambda$. A limitation of using a smaller box size is the inability to estimate the saturation amplitude of fluctuations and angular momentum transport, $\alpha$. Therefore, we do not discuss these issues in this paper. To narrow the dissipation range, the cascade was terminated by fourth-order hyperviscosity and hyperresistivity. The viscous and resistive coefficients are set to the same value (i.e., magnetic Prandtl number Pm is unity). Further details of the numerical setup can be found in Methods.
@@ Results
Figure 1, A and B show the snapshot of the norm of the flow field $\mathbf{u}$ and the magnetic field $\mathbf{B}$ together with typical magnetic field lines, in the planes $x=0$, $y=0$, and $z=0$. Each field is normalized by its root mean square value. The overall structures of both fields are azimuthally elongated due to the Keplerian shear flow. Upon closer inspection, the magnetic field has broad structures, while the flow field is concentrated in relatively smaller patches. The reason for this difference is that the large-scale azimuthal magnetic field $B_{y}$ is preferentially amplified by the $\Omega$ effect  (?), which is evident from the shape of magnetic field lines on the $x=0$ and $z=0$ planes (see also Fig. S1 for the snapshot of all components of $\mathbf{u}$ and $\mathbf{B}$ showing that only $B_{y}$ has intense large-scale structures). One also notices that the magnetic field lines are predominantly azimuthal but have a finite radial component with anticorrelation between $B_{x}$ and $B_{y}$ (consistent with the hybrid PIC simulation of the shearing box magnetorotational turbulence  (?); see also Fig. S1, A and B). Figure 2, A and B show the same snapshot of $\mathbf{u}$ and $\mathbf{B}$ but high-pass filtered by removing the fluctuation with a wavenumber smaller than 20. As we will show later, the fluctuations with the wavenumber greater than $10$ are arguably in the inertial range. The filtered $\mathbf{u}$ and $\mathbf{B}$ exhibit similar morphology, unlike their unfiltered counterparts, suggesting that the cascade tends to be Alfvénic as it moves towards smaller scales. It is also observed that the spatial structures of the filtered fields are elongated along the magnetic field lines. The magnified regions in Fig 2, A and B clearly visualize this elongation. Thus, the large-scale magnetic field effectively acts like a mean field for the fluctuations in the inertial range. One also finds that the azimuthal polarity of the mean magnetic field is not uniform, which is evident from the typical magnetic field lines plotted in Fig. 1 and 2. This is because MRI creates both positive and negative $B_{x}$, which turn into positive and negative $B_{y}$ via shear flow. Therefore, the emergence of the current sheet and magnetic reconnection of a mean field are naturally expected to occur in the simulation, and indeed, we find the chain of multiple plasmoids in the magnified region on the $x=0$ plane in Fig. 2B. Nevertheless, these plasmoids do not fill a substantial portion of the simulation domain but appear only in limited regions.
@@
To explore the properties of small-scale fluctuations, we analyzed the energy spectra of $\mathbf{u}$ and $\mathbf{B}$. Figure 3A and B present the two-dimensional spectrum of magnetic energy. Figure 3A shows the spectrum as a function of $k_{z}$ and $k_{y}$, with the $k_{x}$-direction integrated out, while Fig. 3B displays the spectrum as a function of $k_{z}$ and $k_{x}$, integrating out the $k_{y}$-direction. The snapshot was taken immediately after remapping from the shearing coordinate to the laboratory coordinate, so that the radial wavenumber $k_{x}$ in both coordinates coincides (see Methods for details about the periodic remapping and the time dependency of $k_{x}$). The results indicate anisotropy, specifically $k_{x}\simeq k_{z}>k_{y}$. This aligns with the observation seen in Figs. 1 and 2 that the fluctuations are elongated along the shear and the mean magnetic field on the $x-y$ and $y-z$ planes, while the structures are nearly isotropic on the $x-z$ plane. However, Figure 3C, which displays $k_{x}$ and $k_{z}$ as functions of $k_{z}$ by flattening Figures 3A and B, shows that the anisotropy is scale-independent, unlike the scale-dependent anisotropy commonly seen in other simulations of MHD turbulence with external forcing and mean magnetic field. This is arguably because the mean magnetic field is not exactly $B_{y}$, and the perpendicular component of the wavenumber contaminates $k_{y}$.
@@
Figure 4A shows the omnidirectional spectra of kinetic, magnetic, and total energy. The total energy spectrum is well fitted by a power law of $k^{-5/3}$ when $k\lesssim 100$, where $k$ is the wavenumber, and the spectral slope gets slightly shallower when $k\gtrsim 100$. While the $-5/3$ spectrum was captured in the previously highest-resolution simulation  (?), the shallowing of the slope is found in this study for the first time. We found that the shallowing occurs only in the magnetic field, while the kinetic-energy spectrum is almost completely $k^{-3/2}$ throughout the wavenumber domain. In the previously highest-resolution simulation  (?), the kinetic-energy spectrum was slightly shallower than $k^{-3/2}$. This is likely due to insufficient numerical resolution because the kinetic-energy spectrum becomes shallower than $k^{-3/2}$ in our low-resolution simulation to be shown below.
The shallowing of the magnetic-energy spectrum is not surprising for the following reason. The simulation of MHD turbulence with external forcing indicates that both magnetic- and kinetic-energy spectra converge to $k^{-3/2}$ with the same magnitude  (?). Meanwhile, as the cascade proceeds, the eddy turnover time decreases, causing fluctuations to lose memory of their MRI origin and eventually align with the above prediction  (?, ?). At large scales, magnetic fluctuations are greater than those of kinetic energy and have a steeper spectrum, whereas the kinetic-energy spectrum is already $k^{-3/2}$, as shown in Fig. 4A. Thus, the magnetic-energy spectrum must flatten in order to merge into the kinetic-energy spectrum.
@@
We further investigate the magnetic-energy spectrum by plotting $x,y$, and $z$ components in Fig. 4B. As is consistent with the previous simulation  (?), $B_{y}$ is dominant on the large scale (due to amplification via the $\Omega$ effect) and diminishes quickly in the smaller scales. However, unlike the previous simulation, the $B_{y}$ spectrum becomes shallower at $k\gtrsim 100$, at which the total energy spectrum starts to become shallower than $-5/3$. The second dominant component on the large scale is $B_{x}$, which is presumably due to MRI. Although one might doubt that the shallowing of the spectrum is the numerical roll-up due to the hyper-dissipation, we do not think that this is the case because of the following two reasons. First, only the $B_{y}$ spectrum shows shallowing, while the other fields do not. Second, the shallowing disappears in our low-numerical-resolution simulation, as we will show below. We will further investigate this shallowing of $B_{y}$ later in this paper. Although these two points do not rule out the possibility of numerical roll-up completely, it is currently computationally impossible to validate the presence of shallowing using Laplacian dissipation.
@@
Evidence that our simulation resolved the inertial range is provided in Fig. 4C, which plots the cross-scale energy flux through the wavenumber shell $|\mathbf{k}|=k$. Here, the energy flux $\Pi^{f^{<}}_{g^{>}}(k)$ denotes the energy transfer from the field $f$, where $f=u$ (or $=B$) for the flow (or magnetic) field, with the wavenumber smaller than $k$ to the field $g$ with the wavenumber larger than $k$ (see Methods for the mathematical definition). We find that the total energy flux is fairly constant at $k\gtrsim 10$, which means that the cascade at $k\gtrsim 10$ is in the inertial range. We stress that, in the previous simulations of MRI turbulence, the cross-scale energy flux was not constant  (?), and our simulations is the first time that found the constant energy flux. Note, however, that not only the number of grid points, but the size of the simulation box and the aspect ratio of their simulation are different from ours. We also find that the amount of energy flux “from $u$ and $B$ to $B$” dominates that “from $u$ and $B$ to $u$”, which is consistent with the observation by Ref.  (?).
@@
The nonlinear energy transfer is further investigated in Fig. 5. We define the transfer function $\mathcal{T}_{fg}(Q,K)$ that denotes the energy transfer from the field $f$ in the wavenumber shell $Q\leq|\mathbf{k}|<Q+1$ to the field $g$ in the wavenumber shell $K\leq|\mathbf{k}|<K+1$  (?, ?, ?, ?) (see Methods for the mathematical definition). We find that the dominant energy transfer among all possible combinations of fields is $\mathcal{T}_{BB}$, which is consistent with Ref.  (?). Noticeably, there is a transition at $|\mathbf{k}|\approx 4$. In $|\mathbf{k}|>4$, both $\mathcal{T}_{uu}$ and $\mathcal{T}_{BB}$ are fairly local, and the direction of the cascade is forward. On the other hand, in $|\mathbf{k}|<4$, none of the energy transfers are local. We also find that there is an inverse energy transfer for $\mathcal{T}_{BB}$, suggesting that the large-scale structures of the magnetic field shown in Fig. 1A is formed not just by the $\Omega$ effect but also by the inverse cascade. Furthermore, the transition scale, $|\mathbf{k}|\approx 4$, coincides with the scale where the energy flux becomes constant (Fig. 4C), meaning that the energy cascade in the inertial range is local. Regarding the energy transfers between $u$ and $B$, we find that they are more non-local than the transfer within $u$ or $B$ as $\mathcal{T}_{uB}$ and $\mathcal{T}_{Bu}$ have broad off-diagonal tails. That being said, the non-local transfer from the injection range ($|\mathbf{k}|\leq 4$) peters out as the cascade proceeds. This is evident from Fig. 5, E-H, which shows the contribution of energy transfer from the injection range defined by $\sum_{Q\leq 4}\mathcal{T}_{fg}(K,Q)/\sum_{Q\leq K-1}\mathcal{T}_{fg}(K,Q)$. We find that the contribution of the injection range in $\mathcal{T}_{uu}$ and $\mathcal{T}_{BB}$ disappears immediately below the transition scale $|\mathbf{k}|=4$, which is obvious since these transfers are fairly local, as mentioned above. For $\mathcal{T}_{uB}$ and $\mathcal{T}_{Bu}$, the contribution of the injection range survives down to the relatively smaller scale, but the contribution becomes less than 10% at $|\mathbf{k}|=50$ for $\mathcal{T}_{uB}$ and at $|\mathbf{k}|=40$ for $\mathcal{T}_{Bu}$. This is dramatically different from Ref.  (?), which showed that the box-scale to grid-scale transfer of $\mathcal{T}_{uB}$ was substantial. In short, the two facts that the cross-scale energy flux is constant and that the non-local energy transfer from the injection range peters out manifest that the cascade in our simulation is in the inertial range. This is one of the two main results of this paper.
@@
The other main result of this paper, namely, the partition between slow-magnetosonic and Alfvénic fluctuations in the magnetorotational turbulence, is shown in Fig. 6. In incompressible MHD, the flow and magnetic field of the Alfvén waves are described by $u_{\perp}\hat{\mathbf{k}}_{\|}\times\hat{\mathbf{k}}_{\perp}$ and $\delta B_{\perp}\hat{\mathbf{k}}_{\|}\times\hat{\mathbf{k}}_{\perp}$ while those of the slow-magnetosonic waves are described by $u_{\|}[\hat{\mathbf{k}}_{\|}-(k_{\|}/k_{\perp})\hat{\mathbf{k}}_{\perp}]$ and $\delta B_{\|}[\hat{\mathbf{k}}_{\|}-(k_{\|}/k_{\perp})\hat{\mathbf{k}}_{\perp}]$, respectively, where the hat symbol denotes a unit vector, and the eigenfunction of the slow magnetosonic waves is used. As we found in Figs. 1 and 3, the wavenumber is strongly anisotropic, i.e., $k_{\|}/k_{\perp}\ll 1$. This allows the slow-magnetosonic waves to be represented simply by $u_{\|}$ and $\delta B_{\|}$ (Fig. S3 confirms that neglecting the terms proportional to $k_{\|}/k_{\perp}$ does not change the spectra). Thus, we can decompose the total magnetic and kinetic energy into those of Alfvénic and slow-magnetosonic fluctuations via projection of $\mathbf{u}$ and $\mathbf{B}$ onto the mean magnetic field. However, the global mean magnetic field does not always serve as a mean field for small-scale fluctuations, and thus, we use the method developed by Cho and Lazarian  (?) to decompose the fluctuation and the local mean magnetic field $\mathbf{B}_{0}(\mathbf{r})$. For a given wavenumber $k$, the local mean magnetic field $\mathbf{B}_{0}(\mathbf{r})$ is obtained by filtering the Fourier modes of $\mathbf{B}$ with the wavenumber greater than $k/2$, and the fluctuations are obtained by filtering out the Fourier modes of $\mathbf{u}$ and $\mathbf{B}$ with the wavenumber smaller than $k/2$ or greater than $2k$. Then we decompose the fluctuations of $\mathbf{u}$ and $\mathbf{B}$ into parallel ($u_{\|}$ and $\delta B_{\|}$) and perpendicular ($u_{\perp}$ and $\delta B_{\perp}$) components to $\mathbf{B}_{0}$. Figure 6A shows the spectra of the decomposed fields. First, we confirm the validity of our decomposition by comparing the sum of $u_{\|}$ and $u_{\perp}$ with the total $u$ in Fig. 1A and the sum of $\delta B_{\|}$ and $\delta B_{\perp}$ with the total $B$ in Fig. 1A; the spectra of the total fields are almost perfectly recovered from the sum of the decomposed fields. In terms of magnetic fluctuations, one finds that only the spectrum of slow-magnetosonic fluctuations exhibits flattening at $k\gtrsim 100$, while that of Alfvénic fluctuations has nearly the same spectral index throughout the inertial range. In contrast, the kinetic-energy spectra of both Alfvén and slow-magnetosonic fluctuations are almost perfectly $k^{-3/2}$. Therefore, the flattening of the spectrum seen in Fig. 1A is solely due to the magnetic component of slow-magnetosonic fluctuations. The inset of Fig. 6A shows the same spectra obtained from the simulation with lower resolution $N_{x}\times N_{y}\times N_{z}=512\times 512\times 256$, which manifests that the flattening of the magnetic energy of the slow-magnetosonic waves is absent. Thus, the flattening can be seen only with ultra-high resolution. It can be seen that due to this flattening, the magnetic energy and kinetic energy of slow-magnetosonic fluctuations tend to be equipartitioned, and both spectra approach $k^{-3/2}$. On the other hand, the kinetic and magnetic-energy spectra of Alfvénic fluctuations do not converge in our simulation, and a further higher resolution is required to determine the converged spectral slope. Figure 6B shows the ratio of the energy of slow-magnetosonic fluctuations to that of Alfvénic fluctuations, manifesting that slow-magnetosonic fluctuations have approximately twice stronger energy than the Alfvénic ones. One also finds that the ratio is almost constant throughout the wavenumber domain, suggesting that the coupling between the Alfvénic and slow-magnetosonic fluctuations in MRI turbulence is weak. This is consistent with the recent report on the shearing box simulation which found that the ratio of energy injection between Alfvénic and slow-magnetosonic fluctuations almost equals to the ratio of dissipation which we computed using the reduced MHD  (?).
Finally, we compare our results with the simulation of magnetorotational turbulence solved by the reduced MHD with a near-azimuthal mean magnetic field  (?). Figures 6C and D are reproduced from Ref. (?) and show qualitative consistency with our results in Figs. 6A and B. Specifically, Fig. 6C displays the spectra of the slow-magnetosonic and Alfvénic fluctuations, which are remarkably similar to the spectra at $k\gtrsim 100$ in Figure 6A. Figure 6D shows the ratio between the two fluctuations, with a value of $\approx 2$, exactly matching the result in Fig. 6B. Thus, we conclude that the reduced MHD with a near-azimuthal mean magnetic field effectively captures the features of the inertial range of magnetorotational turbulence.